# TreeKV: Smooth Key-Value Cache Compression with Tree Structures

### v1

@@ Abstract
Efficient key-value (KV) cache compression is critical for scaling transformer-based Large Language Models (LLMs) in long sequences and resource-limited settings. Existing methods evict tokens based on their positions or importance scores, but position-based strategies can miss crucial information outside predefined regions, while those relying on global importance scores resulting in strong regional biases, limiting the KV cache’s overall context retention and potentially impairing the performance of LLMs on complex tasks. Our wavelet analysis reveals that as tokens approach the end of sequence, their contributions to generation gradually increase and tends to diverge more from neighboring tokens, indicating a smooth transition with increasing complexity and variability from distant to nearby context. Motivated by this observation, we propose TreeKV, an intuitive, training-free method that employs a tree structure for smooth cache compression. TreeKV maintains a fixed cache size, allowing LLMs to deliver high-quality output even in long text scenarios. Unlike most compression methods, TreeKV is applicable to both the generation and prefilling stages. It consistently surpasses all baseline models in language modeling tasks on PG19 and OpenWebText2, allowing LLMs trained with short context window to generalize to longer window with a 16x cache reduction. On the Longbench benchmark, TreeKV achieves the best performance with only 6% of the budget at optimal efficiency [^1] We’ll open-source it upon acceptance..
@@ 1 Introduction
Aiming for training-free methods, recent studies propose to cache the KV pairs of only a subset of past tokens, subject to a pre-defined capacity constraint. Those efficient KV cache methods typically address two different stages, prefilling stage and decoding stage. The prefilling stage is primarily the phase where the model develops a representation of the context it receives. The decoding stage focused on generating text output based on the encoded input. Effective management of both stages ensures that the cache operates not just as a temporary storage solution but as a proactive element enhancing overall system efficiency. For decoding stage optimization, some methodsXiao et al. (2023); Han et al. (2024) retain only initial and recent tokens, while othersZhang et al. (2024c); Oren et al. (2024); Liu et al. (2024) select tokens based on their importance scores. Those approaches help manage cache size and allow models to generate sequences longer than the pre-training context length. However, these methods often lack efficient prefilling strategy, leading to high latency due to token-by-token processing. In contrast, there are other methods focus on optimizing the prefilling stage strategy by intelligently populating the cache based on the input context. For example, Li et al. (2024) retained only critical tokens from the context, whileZhang et al. (2024b); Wang et al. (2024) identified key information from KV cache using structural insights.
@@
This work aims to overcome the above limitations. To gain a clearer understanding of the characteristics conveyed by the KV cache, we use wavelet transform to examine the frequency representation of information contributed by tokens at different positions during generation. The results reveal that as tokens approach the end of the sequence, the amplitudes of signals corresponding to different frequency components gradually increase, particularly at higher frequencies. This suggests that the information contributed by a token not only increases, but also becomes more distinct from its neighboring tokens as it approaches the end, indicating a smooth transition with increasing complexity and variability from distant to nearby context. This observation inspired out approach, leading to the design of a structure that is sparse on the left and dense on the right.
@@ 2 Related Work
Another line of research is dedicated to structure-guided long context processing, Zhao et al. (2022) developed a hierarchy for selecting important tokens across layers,He et al. (2024) created a multi-scale tree to efficiently capture the long context dependencies. However, these studies are limited to encoder-based models and cannot be directly applied to pre-trained LLMs without additional tuning, restricting their applicability to generative models.
@@ 3.1 KV Caching
$\mathbf{K}^{(t)}=\begin{pmatrix}\mathbf{K}^{(t-1)}\\ \mathbf{k}^{(t)}\end{pmatrix},\mathbf{V}^{(t)}=\begin{pmatrix}\mathbf{V}^{(t-1)}\\ \mathbf{v}^{(t)}\end{pmatrix}.$
The output of the attention is computed as follows:
$\mathbf{o}^{(t)}=\mathrm{SoftMax}\left(\frac{\mathbf{q}^{(t)}{\mathbf{K}^{(t)}}^{\top}}{\sqrt{d}}\right)\cdot\mathbf{V}^{(t)},$
@@ 3.2 Multi-Level Discrete Wavelet Decomposition
$R(\mathbf{A}_{1},\mathbf{D}_{1})[n]=\begin{cases}\frac{\sqrt{2}}{2}\left(\mathbf{A}_{1}\left[\frac{n+1}{2}\right]+\mathbf{D}_{1}\left[\frac{n+1}{2}\right]\right)&\text{odd }n\\ \frac{\sqrt{2}}{2}\left(\mathbf{A}_{1}\left[\frac{n}{2}\right]-\mathbf{D}_{1}\left[\frac{n}{2}\right]\right)&\text{even }n\end{cases}$
In multi-level discrete wavelet decomposition, the approximation coefficients are further decomposed through repeated single-level decomposition. Specifically, at level $l>1$, the approximation coefficients $\mathbf{A}_{l}[n]$ and the detail coefficients $\mathbf{D}_{l}[n]$ are obtained by filtering and down-sampling $\mathbf{A}_{l-1}[n]$. The coefficients of an $L$-level discrete wavelet decomposition can be organized into a list $[\mathbf{A}_{L},\mathbf{D}_{L},\mathbf{D}_{L-1},\cdots,\mathbf{D}_{1}]$, where $\mathbf{A}_{L}$ gives the lowest frequency coefficient of the signal in the decomposition, and $\mathbf{D}_{L},\cdots,\mathbf{D}_{1}$ give the coefficients at progressively higher frequencies. Applying the single-level reconstruction from high to low levels recovers the signal and the reconstruction on a sole frequency coefficient gives the corresponding frequency component.
@@ 3.3 Observation
$\mathbf{s}={\mathbf{a}^{(t)}}^{\top}\odot{\mathbf{V}^{(t)}}=\begin{pmatrix}\mathbf{a}^{(t)}(1)\mathbf{v}^{(t)}(1)\\ \vdots\\ \mathbf{a}^{(t)}(t)\mathbf{v}^{(t)}(t)\end{pmatrix}$
$\mathbf{a}^{(t)}$ and $\mathbf{V}^{(t)}$ are defined in Section 3.1, we analyze it using multi-level discrete wavelet decomposition defined in Section 3.2. More precisely, we apply the decomposition along the sequence-length dimension, to reconstruct the signal $\text{Rec}(\mathbf{D}_{L})$ of frequency components with only the detail coefficients $\mathbf{D}_{L}$ at level $L$ is,
@@
Then we average the magnitude of each frequency component across all layers, heads, and samples. Figure 2 shows the average magnitude of each frequency component at different positions for a 5-level wavelet decomposition with the Haar wavelet at the 512th generation step across 2000 samples. Since previous work has highlighted the significance of maintaining sinks and recent tokensXiao et al. (2023); Han et al. (2024); Zhang et al. (2024c); Liu et al. (2024); Oren et al. (2024), we exclude the first 32 and the last 32 tokens to concentrate on the middle segment of the sequence in the figure. We observe that at the positions get closer to the end, all frequency components gradually increase, particularly at higher frequencies. This suggests that the information contributed by a token not only increases, but also tends to diverges more from its neighboring tokens as the position approaches the end of sequence, indicating a smooth transition with increasing complexity and variability from distant to nearby context. This insight inspired our method, resulting in a structure designed to be sparse on the left and dense on the right.
@@ 4 TreeKV
Figure 3: Applying TreeKV cache compression method on a 17-length sequence with cache size 4. Each subplot demonstrates the eviction process and state of the cache at the end of step 4, 5, 6 and 17, respectively. Variable idx and selection scope are explained in Section 4.1.
In this section, we introduce our method TreeKV. As the name indicates, the key idea of TreeKV is to organize keys and values in a tree-like hierarchy to enhance smooth cache compression by leveraging temporal locality. Unlike most cache compression methods that focus on either the generation or prefilling phases, TreeKV is applicable to both. We will first outline the compression strategy for the decoding stage, then elaborate its application in the prefilling stage.
@@ 4.1 Decoding Stage
We first define a parameter $c$ to indicate the desired number of key-value pairs to retain in the cache, referred to as cache size. During decoding, the KV pair of new tokens are added sequentially to the cache. Once the cache is at maximum capacity $c$, a tree-based approach strategically evicts KV pair of less important token within a specific eviction scope as generation progresses across layer and head. This eviction scope cycles through the cache from distant to nearby context, ensuring a smooth and balanced distribution of retained tokens. Figure 3 provides an illustration of the cache compression process of TreeKV with a maximum cache size set to 4, each subplot demonstrates the eviction process and the state of the cache at the end of step 4, 5, 6 and 17. The detailed compression procedure is meticulously outlined in Algorithm 1.
@@
Each token is associated with its averaged attention weight (calculated in line 6, 7 and 10 of Algorithm 1). The averaged attention weight reflects the importance score of token in the following generation steps, which serves to prioritize which pairs to retain in the cache. We tested various importance criteria in our experiment, including accumulated and normalized attention weights, and found that averaged attention weight delivers the best performance.
@@
When the maximum cache size $c$ is reached, TreeKV evicts the KV pair of an old token within a “eviction scope” to maintain a fixed cache size. The eviction scope is defined as ($\mathrm{idx}$, $\mathrm{idx+1}$), meaning we remove either the $\mathrm{idx}$-th token or $\mathrm{idx+1}$-th in cache with the lower importance score. This scope cycles through the cache from distant to nearby context based on the scalar $\mathrm{idx}$. For example, in Figure 3, when the eviction process starts at step 5 with $\mathrm{idx}=1$, TreeKV evicts the first token in the scope ($1$, $2$) with the lower importance score, resulting in the removal of “The”. The cache then holds: “sunset”, “slowly”, “over” and “the”. The scaler $\mathrm{idx}$ cycles through 1, 2, 3, and 4, shifting the eviction scope from left to right to remove older tokens first while prioritizing the recent ones. At step 17, upon the arrival of the token “evening,” the $idx$ cycles back to $1$, token “vibrant” from the (“subset”, “vibrant”) is evicted. In the figure, we keep the complete eviction process to make the tree structure visible. Note that our algorithm not just construct trees layer by layer but also facilitates the merging of tokens across different tree levels, maintaining a coherent representation of the input sequence.
@@
1:  Initiate $\mathbf{accum\_attn},\mathbf{count},\mathbf{K}^{(0)},\mathbf{V}^{(0)}$ as empty tensors, scalar $\mathrm{idx}=1$.
@@
7:     $\mathbf{count}=\left(\mathbf{count}\cup\{0\}\right)+\mathbf{1}$.
8:     $\mathbf{accum\_attn}=\left(\mathbf{accum\_attn}\cup\{0\}\right)+\mathbf{a}^{(i)}$.
@@
10:        $\mathbf{avg\_attn}=\mathbf{accum\_attn}/\mathbf{count}$.
11:        if $\mathbf{avg\_attn}_{\mathrm{idx}}>\mathbf{avg\_attn}_{\mathrm{idx+1}}$ then
12:           Evict $\mathrm{(idx-1)}$-th elements in $\mathbf{K}^{(i)}$, $\mathbf{V}^{(i)}$, $\mathbf{count}$, $\mathbf{accum\_attn}$.
@@
Figure 1 (c) shows that TreeKV provides an evenly distributed token map compared to the previous two methods, with a smooth transition from distant coarse-grain to recent fine-grain tokens. Moreover, TreeKV periodically recalibrates the hierarchical structure based on ongoing activity, allowing it to adapt to changes in data dynamically. This adaptability ensures that the cache remains relevant to the current context, enabling the system to consistently provide maximum benefit through efficient storage and retrieval of key-value pairs.
@@ 4.2 Prefilling Stage
Most context compression techniques rely on token-level selectionZhang et al. (2024c); Wang et al. (2024), overlooking the fact that either critical or irrelevant information is often spatially clusteredLi et al. (2024). Selecting individual tokens can compromise contextual integrity and computational speed. Therefore, we adopt a block-level eviction policy for prompt compression tasks, applying our algorithm on blocks instead of tokens. TreeKV first divides the prompt into multiple blocks of size $b$, with each block representing one token in decoding stage. FollowingLi et al. (2024), we use the last block of the prompt as the observation window to query the entire sequence. We obtain the importance score of each block by calculating the averaged importance scores for the $b$ tokens within that block.
@@ 5.1 Setup
Table 3: Performance comparison of H2O, SnapKV and TreeKV on Longbench using LlaMa-3.2-1B as base models. Results are reported with cache sizes of 2048, 4096 and 8192. SnapKV results were obtained using their officially released code, while H2O was implemented by the authors. The best average scores are highlighted in bold. Our model ourperforms H2O and SnapKV on all the cache sizes. The results of Full attention models are shown at the top of the table.

### v2

@@ Abstract
Efficient key-value (KV) cache compression is critical for scaling transformer-based Large Language Models (LLMs) in long sequences and resource-limited settings. Existing methods evict tokens based on their positions or importance scores, but position-based strategies can miss crucial information outside predefined regions, while those relying on global importance scores resulting in strong regional biases, limiting the KV cache’s overall context retention and potentially impairing the performance of LLMs on complex tasks. Our wavelet analysis reveals that as tokens approach the end of sequence, their contributions to generation gradually increase and tends to diverge more from neighboring tokens, indicating a smooth transition with increasing complexity and variability from distant to nearby context. Motivated by this observation, we propose TreeKV, an intuitive, training-free method that employs a tree structure for smooth cache compression. TreeKV maintains a fixed cache size, allowing LLMs to deliver high-quality output even in long text scenarios. Unlike most compression methods, TreeKV is applicable to both the generation and prefilling stages. TreeKV consistently surpasses all baseline models in language modeling tasks on PG19 and OpenWebText2, allowing LLMs trained with short context window to generalize to longer window with a 16x cache reduction. On the Longbench benchmark, TreeKV achieves the best performance with only 6% of the budget at optimal efficiency [^1] We’ll open-source our method upon acceptance..
@@ 1 Introduction
Aiming for training-free methods, recent studies propose to cache the KV pairs of only a subset of past tokens, subject to a pre-defined capacity constraint. Those efficient KV cache methods typically address two different stages, prefilling stage and decoding stage. The prefilling stage is primarily the phase where the model develops a representation of the context it receives. The decoding stage focused on generating text output based on the encoded input. Effective management of both stages ensures that the cache operates not just as a temporary storage solution but as a proactive element enhancing overall system efficiency. For decoding stage optimization, some methods Xiao et al. (2023); Han et al. (2024) retain only initial and recent tokens, while others Zhang et al. (2024c); Oren et al. (2024); Liu et al. (2024) select tokens based on their importance scores. Those approaches help manage cache size and allow models to generate sequences longer than the pre-training context length. However, these methods often lack efficient prefilling strategy, leading to high latency due to token-by-token processing. In contrast, there are other methods focusing on optimizing the prefilling stage strategy by intelligently populating the cache based on the input context. For example, Li et al. (2024) retained only critical tokens from the context, while Zhang et al. (2024b); Wang et al. (2024) identified key information from KV cache using structural insights.
@@
This work aims to overcome the above limitations. To gain a clearer understanding of the characteristics conveyed by the KV cache, we use wavelet transform to examine the frequency representation of information contributed by tokens at different positions during generation. The results reveal that as tokens approach the end of the sequence, the amplitudes of signals corresponding to different frequency components gradually increase, particularly at higher frequencies. This suggests that the information contributed by a token not only increases, but also becomes more distinct from its neighboring tokens as it approaches the end, indicating a smooth transition with increasing complexity and variability from distant to nearby context. This observation inspired our approach, leading to the design of a structure that is sparse on the left and dense on the right.
@@ 2 Related Work
Another line of research is dedicated to structure-guided long context processing. Zhao et al. (2022) developed a hierarchy for selecting important tokens across layer, He et al. (2024) created a multi-scale tree to efficiently capture the long context dependencies. However, these studies are limited to encoder-based models and cannot be directly applied to pre-trained LLMs without additional tuning, restricting their applicability to generative models.
@@ 3.1 KV Caching
$\mathbf{K}^{(t)}=\begin{pmatrix}\mathbf{K}^{(t-1)}\\ \mathbf{k}^{(t)}\end{pmatrix},\quad\mathbf{V}^{(t)}=\begin{pmatrix}\mathbf{V}^{(t-1)}\\ \mathbf{v}^{(t)}\end{pmatrix}.$
The attention scores $\mathbf{a}^{(t)}$ and the output of the attention $\mathbf{o}^{(t)}$ are computed as follows:
$\mathbf{a}^{(t)}=\mathrm{SoftMax}\left(\frac{\mathbf{q}^{(t)}{\mathbf{K}^{(t)}}^{\top}}{\sqrt{d}}\right),\quad\mathbf{o}^{(t)}=\mathbf{a}^{(t)}\cdot\mathbf{V}^{(t)}.$
@@ 3.2 Multi-Level Discrete Wavelet Decomposition
$R(\mathbf{A}_{1},\mathbf{D}_{1})[n]=\begin{cases}\frac{\sqrt{2}}{2}\left(\mathbf{A}_{1}\left[\frac{n+1}{2}\right]+\mathbf{D}_{1}\left[\frac{n+1}{2}\right]\right)&\text{odd }n\\ \frac{\sqrt{2}}{2}\left(\mathbf{A}_{1}\left[\frac{n}{2}\right]-\mathbf{D}_{1}\left[\frac{n}{2}\right]\right)&\text{even }n\end{cases}.$
In multi-level discrete wavelet decomposition, the approximation coefficients are further decomposed through repeated single-level decomposition. Specifically, at level $l>1$, the approximation coefficients $\mathbf{A}_{l}[n]$ and the detail coefficients $\mathbf{D}_{l}[n]$ are obtained by filtering and down-sampling $\mathbf{A}_{l-1}[n]$. The coefficients of an $L$-level discrete wavelet decomposition can be organized into a list $[\mathbf{A}_{L},\mathbf{D}_{L},\mathbf{D}_{L-1},\cdots,\mathbf{D}_{1}]$, where $\mathbf{A}_{L}$ gives the lowest frequency coefficient of the signal in the decomposition, and $\mathbf{D}_{L},\cdots,\mathbf{D}_{1}$ give the coefficients at progressively higher frequencies. Applying the single-level reconstruction from high to low levels recovers the original signal.
@@ 3.3 Observation
$\mathbf{s}={\mathbf{a}^{(t)}}^{\top}\circ{\mathbf{V}^{(t)}}=\begin{pmatrix}\mathbf{a}^{(t)}(1)\mathbf{v}^{(t)}(1)\\ \vdots\\ \mathbf{a}^{(t)}(t)\mathbf{v}^{(t)}(t)\end{pmatrix}.$
$\mathbf{a}^{(t)}$ and $\mathbf{V}^{(t)}$ are defined in Section 3.1, we analyze it using multi-level discrete wavelet decomposition defined in Section 3.2. More precisely, we first apply the multi-level wavelet decomposition along the sequence length dimension to obtain the frequency coefficients at various levels. Next, we reconstruct the signal $\text{Rec}(\mathbf{D}_{L})$, which represents frequency components with only the detail coefficients $\mathbf{D}_{L}$,
@@
Then we average the magnitude of each frequency component across all layers, heads, and samples. Figure 2 shows the average magnitude of each frequency component at different positions for a 5-level wavelet decomposition with the Haar wavelet at the 512th generation step across 2000 samples. Since previous work has highlighted the significance of maintaining sinks and recent tokens Xiao et al. (2023); Han et al. (2024); Zhang et al. (2024c); Liu et al. (2024); Oren et al. (2024), we exclude the first 32 and the last 32 tokens to concentrate on the middle segment of the sequence in the figure.
We observe that as the positions get closer to the end, all frequency components gradually increase, particularly at higher frequencies. This suggests that the information contributed by a token not only increases, but also tends to diverges more from its neighboring tokens as the position approaches the end of sequence, indicating a smooth transition with increasing complexity and variability from distant to nearby context. This insight inspired our method, resulting in a structure designed to be sparse on the left and dense on the right.
@@ 4 TreeKV
Figure 3: Applying TreeKV cache compression method on a 17-length sequence with cache size 4. Each subplot demonstrates the eviction process and state of the cache at the end of step 4, 5, 6 and 17, respectively, while steps 1-4, showing the cache filling process, are omitted. Variable idx and selection scope are explained in Section 4.1.
This section introduces TreeKV, which organizes keys and values in a tree-like hierarchy to enhance smooth cache compression by leveraging temporal locality. Note that, the structure of TreeKV is inspired by our spectral analysis in Sec 3.3. Unlike most cache compression methods that focus on either the generation or prefilling phases, TreeKV is applicable to both. We will first outline the compression strategy for the decoding stage, then elaborate its application in the prefilling stage.
@@ 4.1 Decoding Stage
During decoding, the KV pair of new tokens are added sequentially to the cache. Once the cache is at maximum capacity $c$, a tree-based approach strategically evicts KV pair of less important token within a specific eviction scope as generation progresses across layer and head. The parameter $c$ indicates the desired number of key-value pairs to retain in the cache, referred to as cache size. This eviction scope cycles through the cache from distant to nearby context, ensuring a smooth distribution of retained tokens that is sparse on the left and dense on the right. Figure 3 provides an illustration of the cache compression process of TreeKV with a maximum cache size set to 4, each subplot demonstrates the eviction process and the state of the cache at the end of step 4, 5, 6 and 17, with steps 1-4 representing the cache filling process. The detailed compression procedure is meticulously outlined in Algorithm 1.
@@
Each token is associated with its averaged attention weight $\overline{\mathbf{S}}$ (calculated in line 6, 7, 8 and 10 of Algorithm 1). The averaged attention weight $\overline{\mathbf{S}}$ reflects the importance score of token in the following generation steps, which serves to prioritize which pairs to retain in the cache. We tested various importance criteria in our experiment, including accumulated and normalized attention weights, and found that averaged attention weight delivers the best performance.
@@
When the maximum cache size $c$ is reached, TreeKV evicts an old KV pair within an “eviction scope” to maintain the fixed size. The scaler $\mathrm{idx}$ defines this eviction scope, $\mathrm{idx}$ cycles through the cache from distant to nearby contexts, building a tree structure that is sparse on the left and dense on the right. Specifically, the eviction scope is $\left\{\mathrm{idx},\mathrm{idx+1}\right\}$, indicating that we will remove either the $\mathrm{idx}$-th or the $(\mathrm{idx+1})$-th token in the cache with the lower importance score. For example, in Figure 3, when the eviction process starts at step 5 with $\mathrm{idx}=1$, TreeKV evicts the first token in the scope ($1$, $2$) with the lower importance score, resulting in the removal of “The”. The cache then holds: “sunset”, “slowly”, “over” and “the”. The scaler $\mathrm{idx}$ cycles through 1, 2, 3, and 4, shifting the eviction scope from left to right to remove older tokens first while prioritizing the recent ones. At step 17, upon the arrival of the token “evening,” the $\mathrm{idx}$ cycles back to $1$, token “vibrant” from the (“subset”, “vibrant”) is evicted. In the figure, we keep the complete eviction process to make the tree structure visible. Note that our algorithm not just construct trees layer by layer but also facilitates the merging of tokens across different tree levels, maintaining a coherent representation of the input sequence.
@@
1:  Initiate $\mathbf{S},\mathbf{C},\mathbf{K}^{(0)},\mathbf{V}^{(0)}$ as empty tensors, $\mathrm{idx}=1$.
@@
7:     $\mathbf{C}=\left(\mathbf{C}\cup\{0\}\right)+\mathbf{1}$.
8:     $\mathbf{S}=\left(\mathbf{S}\cup\{0\}\right)+\mathbf{a}^{(i)}$.
@@
10:        $\overline{\mathbf{S}}=\mathbf{S}/\mathbf{C}$. $\triangleright$ Important Scores
11:        if $\overline{\mathbf{S}}_{\mathrm{idx}}>\overline{\mathbf{S}}_{\mathrm{idx}+1}$ then
12:           Evict $(\mathrm{idx}+1)$-th elements in $\mathbf{K}^{(i)}$, $\mathbf{V}^{(i)}$, $\mathbf{C}$, $\mathbf{S}$.
@@
Figure 1 (c) shows that TreeKV provides a more balanced distributed token map compared to the previous two methods, with a smooth transition from distant coarse-grain to recent fine-grain tokens. Moreover, TreeKV updates the hierarchical structure based on a continuous scoring mechanism, allowing it to adapt to changes in data dynamically. This adaptability ensures that the cache remains relevant to the current context, enabling the system to consistently provide maximum benefit through efficient storage.
@@ 4.2 Prefilling Stage
Most context compression techniques rely on token-level selection Zhang et al. (2024c); Wang et al. (2024), overlooking the fact that either critical or irrelevant information is often spatially clustered Li et al. (2024). Selecting individual tokens can compromise contextual integrity and computational speed. Therefore, we adopt a block-level eviction policy for prompt compression tasks, applying our algorithm on blocks instead of tokens. TreeKV first divides the prompt into multiple blocks of size $b$, with each block representing one token in Fig 3. Following Li et al. (2024), we use the last block of the prompt as the observation window to query the entire sequence. We obtain the importance score of each block by calculating the averaged importance scores for the $b$ tokens within that block.
@@ 5.1 Setup
Table 3: Performance comparison of H2O, SnapKV and TreeKV on Longbench using LlaMa-3.2-1B as base models. Results are reported with cache sizes of 2048 and 8192. SnapKV results were obtained using their officially released code, while H2O was implemented by the authors. The best scores are highlighted in bold. Our model outperforms H2O and SnapKV on all the cache sizes. The results of Full attention models are shown at the top of the table.